OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, -3) and has a slope of one half.

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

@iambatman ?

OpenStudy (anonymous):

Y2-y1 over x2-x1

OpenStudy (anonymous):

Sorry y=mx+b

OpenStudy (anonymous):

So This Is The Equation so far 6=1/2 (-3)+b Now You Have To Solve for b

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\]

pooja195 (pooja195):

(x1,y1) m=slope

pooja195 (pooja195):

And where are your options??

OpenStudy (anonymous):

Either equation will give you the same answer

OpenStudy (anonymous):

The choices are y - 6 = 1/2 (3 + x) y = 1/2x - 6 y - 3 = 1/2 (x - 6) x - 2y = 12

OpenStudy (anonymous):

@pooja195 @Niaa08

pooja195 (pooja195):

Ok so we dont need to solve \[ \huge~\rm~y−y_1=m(x−x_1)\] (x1,y1) m=slope Just plug in the given info

OpenStudy (anonymous):

Im not so sure how to, im kinda new at this...

OpenStudy (abb0t):

Answer is \(\sf \color{red}{C}\) in other words, choice number 3.

pooja195 (pooja195):

(6, -3) (x1,y1) x1=6 y1=-3 And the slope is given slope =1/2

OpenStudy (anonymous):

Then what?

pooja195 (pooja195):

plug the info into the formula

OpenStudy (abb0t):

Although, it should be \(\sf \ y+3 = \frac{1}{2} x-6\)

OpenStudy (abb0t):

not -3

OpenStudy (abb0t):

But choice 3 looks like the most correct.

pooja195 (pooja195):

@abb0t please stop giving direct answers and let the user learn

OpenStudy (anonymous):

oops yeah it's +3 not -3

OpenStudy (abb0t):

Yep, then that's the answer @TheDragon :)

pooja195 (pooja195):

Wow.

OpenStudy (anonymous):

Thanks @abb0t @pooja195 for the help

OpenStudy (abb0t):

You're very welcome. If you need anymore just tag me <3